# Sentence Smith: Formally Controllable Text Transformation and its Application to Evaluation of Text Embedding Models

### v2

@@ 4.1.3 Data Instantiation
We use two paraphrase datasets as our base data. The first is PAWS (Yang et al., 2019)666https://huggingface.co/datasets/google-research-datasets/paws-x/tree/main, from which we take the positive pairs. While PAWS provides structurally similar paraphrases—originally designed to challenge paraphrase detection models with adversarial positive and negative pairs—it is not fully ideal for our purpose. Due to the high structural similarity, we anticipate slightly lower difficulty in the challenge pairs generated from this dataset.
To obtain a more comprehensive benchmark, we additionally leverage a more diverse set of ChatGPT-generated paraphrases (Vorobev and Kuznetsov, 2023)777https://huggingface.co/datasets/humarin/chatgpt-paraphrases, henceforth denoted as GPTP. This dataset consists solely of positive paraphrase pairs. Also unlike PAWS, GPTP exhibits greater structural variability, making our resulting challenge set inherently more difficult. We validate this statistically by computing bag-of-words similarity between paraphrase pairs in each dataset: PAWS yields an average similarity of 0.9, whereas GPTP has a much lower average similarity of 0.4 (see Appendix A.2 for details).
@@
After employing our parser , a graph transformation , and the generator on every pair, we end up with triples , where is the original sentence, a paraphrase, and our generated foil (structurally highly similar to , but of another meaning). For post-processing, to maximize the integrity of the generated benchmark, we apply a harsh quality filtering process through our NLI-based validation module: We only retain those “contradiction” sentence pairs with a semantic relationship confidence score greater than 90%. This ensures that the final dataset focuses on maximally validated paraphrase foils. However, we will later also experiment with a different filtering criterion. Final dataset statistics, including distribution of transformations, is shown in Table 1.
@@
4.1.4 Evaluation Measures
Given any text embedding model that maps a pair of texts to a real-valued similarity score, we compute two key evaluation metrics. For triplet accuracy , we consider a set with triplets , where is a text, its paraphrase, and our generated foil that differs from in meaning (but not as much in structure). Formally, we then define triplet accuracy as
where returns 1 if condition is true, and 0 else. As a more standard metric, we also measure the Area Under the Receiver Operator Curve () that represents a softer score directly based on the real valued output scores of embedding models. For this, we dichotomize the dataset into positive pairs and negative pairs , essentially inducing a binary paraphrase classification task. Since, it is hard to say which of the two metrics is more apt for selecting models, models should excel in both and thus in our main results we show averages (Opitz, 2024, Rec. 3 in Section 8) using an harmonic mean.
@@ 4.2.1 Embedding Model Selection
Now with our benchmark fully set up, we are ready to evaluate an off-the shelf text embedding model. We select a range of models that have a good balance of performance and efficiency, with some of them ranking among the top models at the time of writing this paper, and others well known.
Specifically, we evaluate 29 embedding models on our novel datasets, among them quite popular ones like LaBSE Feng et al. (2022), SBERT (Reimers and Gurevych, 2019), Jina (Günther et al., 2023) and E5 (Wang et al., 2022b), others are picked from the MTEB public leaderboard888https://huggingface.co/spaces/mteb/leaderboard according to a performance-efficiency balance.
It is important to note that names like “SBERT” or “E5” represent techniques rather than specific model instances. Therefore, from here on, we reference specific models by their Hugging Face identifiers. For instance, all-mpnet-base-v2 refers to a widely used embedding model that applies the SBERT training methodology to a self-supervised pre-trained MPNET model (Song et al., 2020).
@@ 4.3 Results
Table 3 shows the main result. The top ranked model is sentence-t5-large, showing an AVG score of 0.87, outperforming the worst ranked model SGPT-125M by about 30 percentage points (pp.). On PAWS, several models show strong performance of more than 0.90. This is likely due to the higher initial structural similarity of the paraphrases (for statistics see earlier in §A.4), reducing the effectivness of some foils. On GPTP, with its more varied structural differences, the differences between models grows, and even the best performing model only barely exceeds an AVG of 0.8. For individual metrics on the two datasets, we refer to Appendix A.4: Table 6 shows evaluation on GPTP; Table 7 shows evaluation on PAWS.
@@ 4.4 Analysis
We investigate some additional interesting research questions through our novel datasets.
Manipulation Percentage Underspecification (US) 10.33% Role Swap (RS) 20.84% Antonym Replacement (AR) 30.96% Hypernym Substitution (HS) 8.69% Polarity Negation (PN) 29.17%
Which semantic aspects are most challenging for embedding models?
In Table 4 we show the overall tendency of embedding models to get tricked by a particular foil type. Specifically, we find that negation phenomena are not captured well by the models: Either replacing a word with its antonym (AS), or otherwise changing the factuality of the sentence (PN), an embedding model is not unlikely to be foiled. On the other hand, hypernym substitution is least likely to fail a model, accounting for about 9% of embedding model confusions. This is plausible, since changing a word to its hypernym does cause a meaning change, but it does so more subtly than AS, or PN.
A more fine-grained picture is shown in Figure 4. For each individual embedding model we plot the accuracy across different manipulation types within the GPTP dataset. We make several interesting observations: 1. sentence-t5-large excels in all categories. 2. Many embedding models are moderattely robust to polarity changes (purple bars), an important property for models of which we expect finer linguistic understanding. However, this doesn’t hold true for all models, with a few of them exhibiting a accuracy of equal or below 40%. Notably, the (less recent) all-mpnet-base-v2 performs the most robust across all manipulations, making it seem apt for use-cases where sensitivity across the full breadth of possible linguistic manipulations is paramount.
@@
To ensure the highest faithfulness of our automatically induced challenge cases, we selected only pairs where our NLI-based validation scored high for contradiction. In this experiment, we induce a dataset with subtle pairs and less assured pairs, and thus we fish only those examples that score the neutral label with a probability in between 50% and 80%. Specifically, we are interested how much the final ranking deviates in between the two induced datasets.
The results are shown in Table 8, Appendix A.4. Interestingly, sentence-t5-large, same as in all other experiments before, achieves the top rank, even though the intersection of the two datasets is zero. Overall, the ranking stays mostly stable (Spearmanr correlation coefficient: 0.91), and there are no drastic changes.
@@
Our created testing challenge focuses on linguistic understanding and can be dynamically generated and adapted. On the other hand, static and large benchmarks like MTEB cover a large spectrum of tasks, including STS, but also IR, argument mining, and so on. Therefore, we might expect a different ranking of models. Models that would score high in our benchmark but lower on MTEB could point at unleveraged linguistic strengths of a model that might not be fully reflected by MTEB. On the other hand, a relatively lower ranking in our challenge could signal a ‘red flag’ in a model that could suggest vulnerabilities in difficult cases.
In our Appendix, Table 9, we compare the ranking of our models against the relative rankings on MTEB. Notably, the relatively best ranked model on MTEB (cde-small-v1, actually, as of October 1st, 2024 the best small model under 400M params) only reaches place 20 in our challenge, suggesting poorer fine-level semantic understanding than sentence-t5-large a model of similar size, but based on a sequence-to-sequence pre-trained transformer. Thus, our results point at potential vulnerabilities of cde-small-v1, especially towards negation and polarity. To assess the overall difference from our ranking versus the MTEB ranking, we again calculate the Spearmanr. The result 0.54, suggesting a significant but only moderate alignment of the two benchmarks. While the massively-scaled MTEB contains many datasets, our datasets have the advantage that they contain fresh, unseen benchmarking data that can be dynamically generated, and the models can be diagnosed in linguisticunderstanding.
@@ 5 Conclusion
Our SentenceSmith framework facilitates controlled meaning transformations. We first project a text to an abstract graph-based meaning representation, allowing us to change meaning aspects explicitly (e.g., negation). A generator then produces text from the manipulated graph structure. Finally, we leverage an NLI-based to verify an expected outcome. Manual analysis of the generation quality shows high accuracy, with slightly reduced fluency. We showcased SentenceSmith’s utility in an NLP task: benchmarking and analysis of text embedding models. We used it to produce challenging foils from paraphrase pairs, carefully breaking the paraphrase relation while keeping superficial similarity at maximum. Through this, we shed light on the robustness of text embedding models, as well as their facility to distinguish fine-grained linguistic phenomena.
@@
Our main limitations are directly tied to the model of meaning representation. Indeed, while AMR has the great advantage of large data sets as well as reasonably robust parsers and generators, it also comes with limitations and drawbacks that have been, over the years, carefully outlined by research. There is the mono-linguality (Banarescu et al., 2013), then there may be meaning non-isomorphisms due to ambiguity (Wein, 2025), and lack of some scope (Pustejovsky et al., 2019) and tense aspects (Donatelli et al., 2018). Therefore, it is very important to note that SentenceSmith directly scales with improvements in the area of meaning representations, be it their formal improvement, and further improvement of parsers and generators — indeed, while AMR parsers and generators score high on benchmarks according to a metric (e.g., Vasylenko et al., 2023), these tasks are by far not solved (Manning et al., 2020; Opitz and Frank, 2022a; Yang and Schneider, 2024), as is also supported by our output investigation where we observed some fluency issues. Further, with an eye towards multi-linguality, we see some interesting current developments in cross-lingual parsing in the form of “Universal Representations” (Van Gysel et al., 2021) or highly abstract representations that are taxonomy-aligned (Zhang et al., 2024). Exploiting the wealth of AMR tools, we might effectively use an off-the-shelf MT model to wrap the English parsing and generation process (Uhrig et al., 2021), or use cross-lingual AMR parsing for certain languages (Vanroy and Van de Cruys, 2024; Kang et al., 2024). Given sufficient quality of the MT system, this simple approach could already work for some cross-lingual use-cases of our SentenceSmith.
@@ Acknowledgments
Three authors received funding from the Swiss National Science Foundation (SNSF 213585) and the Luxembourg National Research Fund (17498891) through the Impresso research project.
@@ A.3 Examples of Manipulations
Text Choices Type The majority of Havocs served with the Soviets, but the US and Great Britain also used the planes. T: Most Havocs were utilized by the Soviets, although the US and Great Britain also employed them. F: Havocs also serve with the Soviet Union but the majority of their use is by the US and Britain. RS Universities minister David Willetts said all universities can do is ask students if they have booked a flight home. T: According to David Willetts, universities can only inquire about students’ flight reservations for their return home. F: University Minister David Blankts said the university could no longer do the asking whether students had booked flights home. AR It was known as a relatively easy plane to fly with good handling during takeoff and landing T: The aircraft was recognized for its ease of operation and smooth handling during takeoff and landing F: It is known that it is not a relatively easy flying plane that is well-handled when taking off and landing. PN The reason for the crash is not known T: The cause of the accident is uncertain. F: The reason for the crash was known US carcasses were also found near the contaminated watering holes. T: The polluted watering holes were also discovered to have nearby carcasses. F: The carcass was also found far from the contaminated water hole. AR The auction house also sold one of the two gun belts owned by Jesse James at the time of his death. T: One of Jesse James’ two gun belts at the time of his death was also sold by the auction house. F: The auction house also gave up one of Jesse James’s two gun belts when he died. HS Jean-Marc Wenger, who lives in Klingau, found the gold. T: The gold was discovered by Jean-Marc Wenger, a resident of Klingau. F: Jean-Marc Klonau a Wenger resident found gold RS Jean-Marc Wenger, who lives in Klingau, found the gold. T: The gold was discovered by Jean-Marc Wenger, a resident of Klingau. F: Jean-Marc Wenger who lives in Klanau has lost gold. AR But at the moment it is a complete mystery. T: Currently, it remains an enigma. F: But it is not a complete mystery at the moment. PN Thibaut Courtois was out quickly to thwart Sterling as Liverpool looked to get back in the game. T: Liverpool attempted to make a comeback, but Thibaut Courtois swiftly prevented Sterling from scoring F: When Liverpool looked to get back in the game Thibaut Sterling was quick out to thwart Courttois. RS Willis visited the Neon Museum in 2013 to celebrate her 90th birthday. T: In 2013, Willis marked her 90th birthday by visiting the Neon Museum. F: In 2013 to celebrate his 90th birthday he visited the Neo Museum. US AT least 3,000 Brussels bureaucrats earn more than David Cameron, it emerged yesterday. T: Yesterday, it was revealed that over 3,000 Brussels bureaucrats earn a higher salary than David Cameron. F: It emerged yesterday that at least 3000 bureaucrats in Brussels earn less than David Cameron. AR We were going in water until hit the hill and spun. T: We were traveling through water until we hit the hill and spun. F: We went up the hill until the water hit and we spun. RS But you need to get somebody like Warren to do it. T: You should find someone similar to Warren to handle it. F: But you don’t need to get somebody like Warren to do it. PN Julian E. Zelizer says Democrats should be questioning themselves on several key points. T: Democrats ought to be reflecting on various crucial aspects, according to Julian E. Zelizer. F: Julian E. Zelizer said Democrats should question him on several key points. RS

### v4

@@ 4.1.2 Data Instantiation
We use two paraphrase datasets as our base data. The first is PAWS (yang-etal-2019-paws)666https://huggingface.co/datasets/google-research-datasets/paws-x/tree/main, from which we take the positive pairs. While PAWS provides structurally similar paraphrases—originally designed to challenge paraphrase detection models with adversarial positive and negative pairs—it is not fully ideal for our purpose. Due to the high structural similarity, we anticipate slightly lower difficulty in the challenge pairs generated from this dataset.
To obtain a more comprehensive benchmark, we additionally leverage a more diverse set of ChatGPT-generated paraphrases (chatgpt_paraphrases_dataset)777https://huggingface.co/datasets/humarin/chatgpt-paraphrases, henceforth denoted as GPTP. This dataset consists solely of positive paraphrase pairs. Also unlike PAWS, GPTP exhibits greater structural variability, making our resulting challenge set inherently more difficult. We validate this statistically by computing bag-of-words similarity between paraphrase pairs in each dataset: PAWS yields an average similarity of 0.9, whereas GPTP has a much lower average similarity of 0.4 (see Appendix A.2 for details).
@@
After employing our parser , a graph transformation , and the generator on every pair, we end up with triples , where is the original sentence, a paraphrase, and our generated foil (structurally highly similar to , but of another meaning). For post-processing, to maximize the integrity of the generated benchmark, we apply a harsh quality filtering process through our NLI-based validation module: We only retain those “contradiction” sentence pairs with a semantic relationship confidence score greater than 90%. This ensures that the final dataset focuses on maximally validated paraphrase foils. However, we will later also experiment with a different filtering criterion. Final dataset statistics, including distribution of transformations, are shown in Table 1.
@@ 4.2.1 Embedding Model Selection
With our benchmark fully set up, we evaluate 29 off-the-shelf text embedding models. The selection aims to balance performance and efficiency, comprising top-ranked models at the time of writing as well as other widely used baselines. These include LaBSE feng-etal-2022-languagelabse, SBERT (reimers-gurevych-2019-sentence), Jina (gunther-etal-2023-jina), and E5 (wang2022texte5). Additional models were drawn from the MTEB leaderboard888https://huggingface.co/spaces/mteb/leaderboard to provide broader coverage of model types.
It is important to note that names like “SBERT” or “E5” represent techniques rather than specific model instances. Therefore, from here on, we reference specific models by their Hugging Face identifiers. For instance, all-mpnet-base-v2 refers to a widely used embedding model that applies the SBERT training methodology to a self-supervised pre-trained MPNET model (NEURIPS2020_c3a690bempnet).
4.2.2 Evaluation Measures
Consider any model that maps a pair of texts to a real-valued “similarity score” (in our case we have text embedding models that construct the embeddings, and the cosine similarity builds the similarity). To assess the quality of such a model on our induced benchmark, we compute two key evaluation metrics. The first is triplet accuracy , a simple and interpretable classification-oriented score that measures the ratio of cases where the embedding model remains robust to the foil. We consider a set with triplets , where is a text, its paraphrase, and our generated foil designed to mislead the model into assigning it a higher similarity score. Then
where returns 1 if the condition is true, and 0 else. As a more standard metric that is commonly used to evaluate tasks where a floating value (here: similarity) must be compared against a discrete label (here: paraphrase/not-paraphrase), we also measure the Area Under the Receiver Operator Curve () that represents a softer score directly based on the real valued output scores of embedding models. For this, we dichotomize the dataset into positive pairs and negative pairs , essentially inducing a binary paraphrase classification task. Since it is hard to say which of the two metrics is generally more informative, models should clearly excel in both and thus (10.1162/tacl_a_00675, Rec. 3 in Section 8), to compute a single “performance” number for a given dataset, we are using an harmonic mean of TACC and AUC.
@@ 4.3 Results
Table 3 shows the main results. The top-ranked model is sentence-t5-large, showing an AVG score of 0.87, outperforming the worst-ranked model SGPT-125M by about 30 percentage points (pp.). On PAWS, several models show strong performance of more than 0.90. This is likely due to the higher initial structural similarity of the paraphrases (for statistics, see the Appendix, §A.4), reducing the effectiveness of some foils. On GPTP, with its more varied structural differences, the differences between models grows, and even the best performing model only barely exceeds an AVG of 0.8. For individual metrics on each dataset, see Appendix A.4: Table 6 shows evaluation on GPTP, and Table 7 on PAWS.
@@ 4.4 Analysis
A fine-grained picture is shown in Figure 4. For each embedding model, we plot the accuracy across different manipulation types within the GPTP dataset. We make several interesting observations: 1. sentence-t5-large excels in all categories. 2. Many embedding models are moderately robust to polarity changes (purple and red bars), an important property for models of which we expect finer linguistic understanding. However, this is not consistent across models, with some exhibiting a accuracies at or below 40%. Notably, the (less recent) all-mpnet-base-v2 shows the most balanced performance across all manipulations, suggesting its suitability for use cases requiring uniform sensitivity to diverse linguistic manipulations. On average (AVG, right column), role swap is the most difficult transformation, while hypernym substitution is the easiest, though the differences are not pronounced. Overall, the results indicate considerable potential for improvement in embedding models’ fine-grained linguistic understanding.
@@
To ensure the highest faithfulness of our automatically induced challenge cases, we selected only pairs where our NLI-based validation assigned a high probability of contradiction. In this experiment, we construct an alternative dataset containing more ambiguous cases by selecting pairs labeled as neutral with probabilities between 50% and 80%. Specifically, we examine how much the final ranking deviates between the two datasets.
The results are shown in Table 8, Appendix A.4. Interestingly, sentence-t5-large, consistent with all previous experiments, again achieves the top rank, even though the intersection of the two datasets is zero. Overall, the ranking stays mostly stable (Spearman’s rank correlation coefficient: 0.91), with no substantial changes.
@@
Our created testing challenge focuses on linguistic understanding and can be dynamically generated and adapted. On the other hand, static and large benchmarks like MTEB cover a large spectrum of tasks, including STS, but also IR, argument mining, and so on. Therefore, we might expect a different ranking of models. Models that would score high in our benchmark but lower on MTEB could point at unleveraged linguistic strengths of a model that might not be fully reflected by MTEB. On the other hand, a relatively lower ranking in our challenge may indicate a model weakness and suggest vulnerabilities in difficult cases.
In our Appendix, Table 9, we compare the ranking of our models against the relative rankings on MTEB. Notably, the relatively best ranked model on MTEB (cde-small-v1, as of October 1st, 2024, the best model of a size smaller than 400M params) only reaches place 20 in our challenge, suggesting poorer semantic sentence understanding than sentence-t5-large, a model of similar size based on a seq-to-seq T5 transformer. Thus, our results suggest potential vulnerabilities of cde-small-v1, especially towards negation and polarity. To assess the overall difference between our ranking versus the MTEB ranking, we again calculate the Spearman’s rank correlation. We receive a score of 0.54, suggesting a significant but only moderate alignment of the two benchmarks. While MTEB contains a large number of datasets, constituting a very large overall benchmark, our datasets have the advantage that they contain fresh, unseen benchmarking data that can be dynamically generated, and that allow for targeted probing of models’ linguistic understanding.
Embeddings from very large models.
We additionally use our data to probe sentence understanding of text embeddings from two very large language models: E5 with MISTRAL LLM 7B as backbone (wang2022texte5; Jiang2023Mistral7), and GTE with QWEN 7B LLM as backbone li2023towards; bai2023qwen.
Metric LOWER E5-MISTRAL GTE-QWEN UPPER TACC 0.326 0.416 0.540 0.820
Interestingly, the resulting scores (Table 4) are surprisingly low, placing the models in the lower-middle rankings.999The weakest aspect is polarity, where both models get more than 50% of cases wrong.. In contrast, smaller models, specifically T5-based embeddings, perform better. Our hypothesis is that, since these LLM-based models are optimized for retrieval and longer-context tasks, their embeddings may be less sensitive to fine-grained aspects of sentence meaning.
@@ 5 Conclusion
Our SentenceSmith framework facilitates controlled and highly transparent meaning transformations of text. We showcased SentenceSmith’s utility in an NLP task: Dynamic benchmarking and analysis of text embedding models. Concretely, we used it to produce challenging foils from paraphrase pairs, carefully breaking the paraphrase relation while a high degree of superficial similarity. Through this, we shed light on the robustness of text embedding models and their ability to distinguish fine-grained linguistic phenomena.
@@
We note two main types of limitations: those that relate to the model of meaning representation, and those that affect conclusions drawn from our automatically induced benchmark.
Meaning Representation Model.
It is important to note that SentenceSmith directly scales with improvements in the area of meaning representations. While AMR has the great advantage of large data sets as well as reasonably robust parsers and generators, it also comes with limitations and drawbacks that have been, over the years, carefully outlined by research. There is the mono-linguality (banarescu2013abstract), then there may be meaning non-isomorphisms due to ambiguity (wein-2025-ambiguity), and lack of some scope (pustejovsky-etal-2019-modeling) and tense aspects (donatelli-etal-2018-annotation). Moreover, while AMR parsers and generators seem to score high on benchmarks (e.g., vasylenko-etal-2023-incorporating), these tasks remain far from solved (manning-etal-2020-human; opitz-frank-2022-better; yang-schneider-2024-relative), as is also supported by our output investigation where we observed some fluency issues, that may both be due to generation or parsing issues. This setup restricts our framework to English, because robust parsers and generators are not yet widely available for other languages. However, in the context of multilinguality, promising advances in cross-lingual parsing have emerged through “Universal Representations” (van2021designing). Exploiting the wealth of AMR tools, we might also effectively use an off-the-shelf MT model to wrap the English AMR parsing and generation process (uhrig-etal-2021-translate), or use cross-lingual AMR parsing for certain languages (vanroy-van-de-cruys-2024-less; kang-etal-2024-cross-lingual). Given sufficient quality of the MT system, this simple approach could already work for some cross-lingual use-cases of our SentenceSmith.
On drawing conclusions about embedding model performance.
Embedding models are widely used in research and industry, making it important to investigate differences in their performance. Our application study and its induced dataset do allow us to test models’ sensitivity to paraphrases while also highlighting the specific linguistic phenomena that drive this sensitivity. However, different embedding objectives and downstream settings may prioritize relevance as the primary goal, in which case this testing strategy may not provide the most informative signal for model selection. The ranking of models in our setting is conditional on their “similarity” scores serving as proxies for semantic similarity, rather than on a neutral or universal evaluation of embedding quality. Thus, the results should not be misinterpreted as a general quality difference between embedding models. Instead, they reflect the constraint that performance (or, more cautiously, behavior) is evaluated only with respect to specific linguistic paraphrase phenomena. A strength of this targeted evaluation, however, is that it enables us to diagnose model weaknesses in particular categories. Future work may adapt the manipulation rules in ways so that additional objectives can be effectively evaluated.
@@ Acknowledgments
We thank the anonymous reviewers for their comments, and the meta-reviewer for their detailed feedback. Three authors received funding through the project Impresso – Media Monitoring of the Past II. Beyond Borders: Connecting Historical Newspapers and Radio. Impresso is a research project funded by the Swiss National Science Foundation (SNSF 213585) and the Luxembourg National Research Fund (17498891).
@@ A.3 Examples of Manipulations
| Text | Choices | Type |
| The majority of Havocs served with the Soviets, but the US and Great Britain also used the planes. | T: Most Havocs were utilized by the Soviets, although the US and Great Britain also employed them. F: Havocs also serve with the Soviet Union but the majority of their use is by the US and Britain. | RS |
| Universities minister David Willetts said all universities can do is ask students if they have booked a flight home. | T: According to David Willetts, universities can only inquire about students’ flight reservations for their return home. F: University Minister David Blankts said the university could no longer do the asking whether students had booked flights home. | AR |
| It was known as a relatively easy plane to fly with good handling during takeoff and landing. | T: The aircraft was recognized for its ease of operation and smooth handling during takeoff and landing. F: It is known that it is not a relatively easy flying plane that is well-handled when taking off and landing. | PN |
| The reason for the crash is not known. | T: The cause of the accident is uncertain. F: The reason for the crash was known. | US |
| Carcasses were also found near the contaminated watering holes. | T: The polluted watering holes were also discovered to have nearby carcasses. F: The carcass was also found far from the contaminated water hole. | AR |
| The auction house also sold one of the two gun belts owned by Jesse James at the time of his death. | T: One of Jesse James’ two gun belts at the time of his death was also sold by the auction house. F: The auction house also gave up one of Jesse James’s two gun belts when he died. | HS |
| Jean-Marc Wenger, who lives in Klingau, found the gold. | T: The gold was discovered by Jean-Marc Wenger, a resident of Klingau. F: Jean-Marc Klonau a Wenger resident found gold. | RS |
| Jean-Marc Wenger, who lives in Klingau, found the gold. | T: The gold was discovered by Jean-Marc Wenger, a resident of Klingau. F: Jean-Marc Wenger who lives in Klanau has lost gold. | AR |
| But at the moment it is a complete mystery. | T: Currently, it remains an enigma. F: But it is not a complete mystery at the moment. | PN |
| Thibaut Courtois was out quickly to thwart Sterling as Liverpool looked to get back in the game. | T: Liverpool attempted to make a comeback, but Thibaut Courtois swiftly prevented Sterling from scoring. F: When Liverpool looked to get back in the game Thibaut Sterling was quick out to thwart Courttois. | RS |
| Willis visited the Neon Museum in 2013 to celebrate her 90th birthday. | T: In 2013, Willis marked her 90th birthday by visiting the Neon Museum. F: In 2013 to celebrate his 90th birthday he visited the Neo Museum. | US |
| At least 3,000 Brussels bureaucrats earn more than David Cameron, it emerged yesterday. | T: Yesterday, it was revealed that over 3,000 Brussels bureaucrats earn a higher salary than David Cameron. F: It emerged yesterday that at least 3000 bureaucrats in Brussels earn less than David Cameron. | AR |
| We were going in water until hit the hill and spun. | T: We were traveling through water until we hit the hill and spun. F: We went up the hill until the water hit and we spun. | RS |
| But you need to get somebody like Warren to do it. | T: You should find someone similar to Warren to handle it. F: But you don’t need to get somebody like Warren to do it. | PN |
| Julian E. Zelizer says Democrats should be questioning themselves on several key points. | T: Democrats ought to be reflecting on various crucial aspects, according to Julian E. Zelizer. F: Julian E. Zelizer said Democrats should question him on several key points. | RS |